[clas_members] [New Logentry] RC Daily Summary August 17, 2022

yprok at jlab.org yprok at jlab.org
Wed Aug 17 10:55:02 EDT 2022

Logentry Text:
Tuesday, August 16

Day Shift:  Took data on carbon until the start of pass change at 9am. Hall went into controlled access, and target group replaced C with the NH3 target. Target polarization reached 52%.

Swing Shift: Hall was in beam permit, and we prepared to tune beam to the tagger, but beam never came. 
At 6pm,  PD informed us that there was a major repair needed at the accelerator, and it will be delayed until next morning.  We ramped down the magnets, and canceled the rest of the shift. 

Wednesday, August 17th

Owl shift:  Shift canceled. 

Day Shift: Accelerator is working on 1L12 transformer repairs.  We are not expecting beam in the hall until noon. 

We will have the daily RC meeting today at 2:15 PM:

RC Meeting Coordinates:
Meeting ID: 160 867 4253
Passcode: 786506

Data taking progress:  No data has been taken since 9am of 08/16/2022


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4030130
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220817/1286fdcb/attachment.html>

More information about the clas_members mailing list